OpenStudy (anonymous):

Solve the system of equations. 4x + 7y = 6 and 6x + 5y = 20

OpenStudy (anonymous):

10x+12y=26 10+12= 26

OpenStudy (anonymous):

4x + 7y = 6 and 6x + 5y = 20 equal 24x+ 42y= 36 24x+ 20y= 80 together, you get 22y= -44 y= -2 4x+ 7(-2) =6 4x= -20 x= -5